15x-10y=1
15y=2+5x
whats the solution?

Answer by scott8148(6628)

15x = 10y + 1

multiplying 2nd eqn by 3 ___ 45y = 6 + 15x

substituting ___ 45y = 6 + 10y + 1

35y = 7 ___ y = 1/5

substituting ___ 15x = 10(1/5) + 1

15x = 3 ___ x = 1/5